OpenStudy (anonymous):

Let f(x) = sin(x) cot(x). (a) Use the product rule to find f ' (x). (b) True or false: for all real numbers x, f(x) = cos(x). (c) Explain why the function that you found in (a) is almost the opposite of the sine function, but not quite. (Hint: convert all of the trigonometric functions in (a) to sines and cosines, and work to simplify. Think carefully about the domain of f and the domain of f ' .)

OpenStudy (kc_kennylau):

Where are you stuck at?

OpenStudy (anonymous):

I can not solve it at all

OpenStudy (kc_kennylau):

Do you know what the product rule is?

OpenStudy (anonymous):

yes I know

OpenStudy (anonymous):

pq'+p'q

OpenStudy (kc_kennylau):

Then substitute values inside

OpenStudy (anonymous):

here is the problem. I don't know how

OpenStudy (kc_kennylau):

p=sin(x) q=cot(x)

OpenStudy (kc_kennylau):

p'=cos(x) q'=-csc^2(x)